OpenStudy (anonymous):

cosθ-11sinθ=6/cosθ [0°≤θ≤360°] \(\cos ^{2} θ-11\sin θ\cos θ=6\) \(1-\sin ^{2}θ-11\sinθ\cosθ=6(\sin ^{2}+\cos ^{2}θ)\) and then i don't know how to solve.

OpenStudy (john_es):

May be with a change of variable, \[z=\sin\theta\] \[1-z^2-11z\sqrt{1-z^2}=6\Rightarrow-11z\sqrt{1-z^2}=5+z^2\] And solving for z.

OpenStudy (anonymous):

my teacher gave me i tip that the second step is need to be continued

OpenStudy (anonymous):

we didn't study π with trigonometric

OpenStudy (john_es):

Ok, the trick is completing squares.

OpenStudy (callisto):

\[cosθ-11sinθ=\frac{6}{cos\theta}\]\[cos^2\theta - 11sin\theta \cos\theta = 6\]\[cos^2\theta - 11sin\theta \cos\theta = 6(sin^2\theta+cos^2\theta)\]\[cos^2\theta - 11sin\theta \cos\theta - 6sin^2\theta-6cos^2\theta=0\]\[5cos^2\theta + 11sin\theta \cos\theta+ 6sin^2\theta=0\]Factorize it. \[(5\cos\theta+6\sin\theta)(cos\theta+\sin\theta)=0\] I guess you can solve it now

OpenStudy (john_es):

Yes. :)

OpenStudy (callisto):

A very nice hint from your teacher :)

OpenStudy (anonymous):

@Callisto thank you, i want to ask is there any program in the calculator fx-50FH of factorizing? i hate factorizing very much ..

OpenStudy (callisto):

You can use the quadratic formula to cheat... But I usually do it by hand. Practice makes perfect :)

OpenStudy (anonymous):

quadratic formula can do this type of question? tests don't have enough time for me to factorize my hand haha

OpenStudy (callisto):